22) Consider a square which is 1.0 m on a side. Charges are placed at the corners of the square as follows: +4.0 Н.С at (0, 0); +4.0 μC at (1, 1); +3.0 HC at (1,0);-30 μC at (0, 1). What is the magnitude of the electric field at the squares center? A) 1.7 x105 N/C B) 1.3 x105 N/C C)1.5 105 N/C D) 1.9 x 105 N/C E1.1 105 N/C
Answer #1

Here +4uC and +4uC will cancel effect of each other as both are diagonal opposite and hence electric field at center will be due to +3 uC and -3uC charge

Now E = kQ1/r^2 +kQ2/r^2

Here r = 1/sqroot2 so r^2 =1/2

So E = 3×10^-6×9×10^9/(1/2) +3×10^-6×9×10^9/(1/2)

E = 2×(3×10^-6)×9 ×10^9/(1/2) = 1.08×10^5 N/C

Or E = 1.1 × 10^5 N/C

Hence option E is correct